OpenStudy (anonymous):

What is the symmetric function for \(α^5+β^5\)?

OpenStudy (anonymous):

Where \(\alpha\) and \(\beta\) are roots of a quadratic polynomial.

OpenStudy (sasogeek):

I believe it is the expansion of that and then substituting the values for \(\alpha + \beta \) and \(\alpha\beta\)

OpenStudy (anonymous):

I don't know, I have no idea about this.

OpenStudy (anonymous):

I don't even know what question means. If someone could explain me the question.

OpenStudy (amistre64):

i cant make any sense of what google has to offer in this

OpenStudy (anonymous):

All quadratic functions are symmetric, right? \[(x-a)(x-b)=0\] \(a\) and \(b\) are the roots and that's all I can make out of it. Earlier FoolForMath posted a definition "A function of α and β is said to be symmetric if it remains unchanged when α and β are interchanged".

OpenStudy (akshay_budhkar):

but it is not asking for symmetric function of the quadratic equation

OpenStudy (anonymous):

Symmetric function of roots: A functions of α and β is said to be symmetric if it's remain unchanged when α and β are interchanged. \[ α^5+β^5 = (α^3+β^3)(α^2+β^2)-α^2β^2(α+β) \]

OpenStudy (anonymous):

Unrelated Fun: http://www.damnlol.com/i/7355afe7994b2baf4125c449794140c6.jpg

OpenStudy (anonymous):

But what does this Graphically imply? And what is \(\alpha^5+\beta^5\)? I can a plot a function and show \(\alpha\) and \(\beta\) but I don't a thing what \(\alpha^5+\beta^5\) or \(|\alpha + \beta|\) is.

OpenStudy (anonymous):

"...I don't know* a thing..."

OpenStudy (anonymous):

I am not aware of any graphical significance but I have used it some problems where a quadratic equation is given and we need find these ...

OpenStudy (anonymous):

Oh :-( then I guess I would remain confused about this concept

OpenStudy (anonymous):

btw FoolForMath what is the function? \(↓ ↓ \) This is just a algebraic manipulation, right? This can't be a function?! \[α^5+β^5 = (α^3+β^3)(α^2+β^2)-α^2β^2(α+β)\]

OpenStudy (anonymous):

What do you after the manipulation, do you replace the \(\alpha\) and \(\beta\) using sums and products property of a quadratic expression?

OpenStudy (sasogeek):

why does amistre have to mention google in his answers lol, is it some fun fashion kinda thing with respect to answering questions?

OpenStudy (anonymous):

But even if you replace the α and β using sums and products property of a quadratic expression, you won't get a function. You will only get a constant value. So what is the function?

OpenStudy (sasogeek):

this is quite messy lol

OpenStudy (mr.math):

I don't really understand the question, could you explain?

OpenStudy (anonymous):

Yeah, at least for me it is. @saso Me neither Mr.Math, FoolForMath asked this question [at] http://openstudy.com/users/sasogeek#/updates/4f3f4c0ae4b0534c53cbbc1d

OpenStudy (mr.math):

The question doesn't make sense to me. We can write a function \(f(\alpha,\beta)=\alpha^5+\beta^5.\) This function is symmetric since \(f(\alpha,\beta)=f(\beta,\alpha) \) for all \(\alpha\) and \(\beta\) in the domain of \(f\). But I don't see this answers the question that I don't understand :-D

OpenStudy (anonymous):

But αand β are constants, why would they ever make a function?

OpenStudy (mr.math):

Oh.

OpenStudy (anonymous):

Where α and β are roots of a quadratic polynomial.

OpenStudy (anonymous):

I can't find my question paper, else I would have posted one example problem.

OpenStudy (anonymous):

Maybe I am overthinking this :-/

OpenStudy (anonymous):

yes you are.

OpenStudy (mr.math):

Yes, I think there's something missing in the question.